OpenStudy (anonymous):

How do I solve 8cos x = (sec x)^2?

OpenStudy (anonymous):

sec^2x = 1/cos^2x

OpenStudy (anonymous):

8cos^3x=0

OpenStudy (anonymous):

cos(pi/2) = 0

OpenStudy (anonymous):

so 8cos^3(pi/2)=0

OpenStudy (anonymous):

I got lost on the first step.. Sorry, can you explain it more?

OpenStudy (anonymous):

sure, first you need to memorize all of the identitites memorize : sec(x) = 1/cos(x)

OpenStudy (anonymous):

whoa whoa whoa, \[8\cos^2(x) = 1\] not zero...

OpenStudy (anonymous):

so sec^2(x) = 1/cos^2(x)

OpenStudy (anonymous):

ah right 1 my bad

OpenStudy (anonymous):

 So like this?